# Julian Martinez-Iriarte, and Gabriel Montes-Rojas

### v3

@@ 3 Identification: a control function approach
Control Function Condition: there exists $\tilde{V}$ which satisfies $E[\varepsilon|D,X,Z]=E[\varepsilon|\tilde{V}]=r(\tilde{V})$.
Assumption 3 is a standard control function condition, see, e.g., Blundell and Powell (2008) and Florens, Heckman, Meghir, and Vytlacil (2008). Under this condition, the conditional expectation of the response variable in equation (1) can be written as
@@
$=D\alpha_{1}+X^{\prime}\alpha_{2}+g(D,X)r(\tilde{V}).$ (4)
It should be noted that, from the previous display, the skedastic term in the structural equation makes even the standard general identification procedures based on the control function fail. Further assumptions are required to identify $\alpha_{1}$. To illustrate this, we use Example 1 and follow the general identification method for nonparametric triangular models in Newey, Powell, and Vella (1999, Theorem 2.3). For simplicity, we assume that there are no additional exogenous covariates, $X$.
@@
Consider the model in Example 1 and Assumption 3. Define
$f(D,Z):=E[Y|D,Z]$ $=D\alpha_{1}+g(D)E[\varepsilon|D,Z]$
$=D\alpha_{1}+g(D)r(\tilde{V})$
$=D\alpha_{1}+g(D)r(\psi(D,Z)),$
where $\tilde{V}:=\psi(D,Z)$ is an assumed identified control function from the first-stage. Now consider the following derivatives
$\frac{\partial f}{\partial D}$ $=\alpha_{1}+g_{D}r+gr_{\tilde{V}}\psi_{D}$
$\frac{\partial f}{\partial Z}$ $=gr_{\tilde{V}}\psi_{Z},$
@@
Note, however, that one cannot identify $\alpha_{1}$ because of the additional term $g_{D}r$, that is,
@@
The previous example shows that further assumptions are required for identification of the parameter $\alpha_{1}$. The Assumption 4 below imposes a polynomial function on the skedastic function. This condition is the same as in Florens, Heckman, Meghir, and Vytlacil (2008), except that it includes additional exogenous covariates.
@@
Here $X^{\prime k_{g}}$ refers to the element-by-element $k_{g}$ power of $X^{\prime}$. This could be replaced by any other polynomial that may include interactions. Finally, to make the model tractable, we impose the following polynomial structure on the CF. This simplifies identification, estimation, and inference.
@@
Polynomial Control Function: $r(\tilde{V})$ is a polynomial of degree $k_{v}$ in $\tilde{V}$ that does not contain a constant. That is $r(\tilde{V})=\sum_{j=1}^{k_{v}}\mu_{j}\tilde{V}^{j}$.
An important condition in Assumption 5 is the exclusion of a constant term in the function $r(\tilde{V})$. Otherwise, we may allow for $r(\tilde{V})$ to include a constant, but restrict $g(D,X)$ to not contain a linear term with $D$ and $X$. The intuition for these restrictions is simple. Suppose there is only one endogenous regression $D$ and the both the skedastic function and the CF contain a constant as: $g(D)=1+\delta D$ and $r(\tilde{V})=1+\tilde{V}$. Then, $g(D)r(\tilde{V})=1+\tilde{V}+\delta D+\delta D\tilde{V}$, and in this case, one is not able to identify $\alpha_{1}$ because of the term $\delta D$. This is also highlighted by Florens, Heckman, Meghir, and Vytlacil (2008).
The next lemma considers the identification of $\alpha_{1}$. It shows that the parameter of interest is identified by a regression of $Y$ on transformations of $(D,X,\tilde{V})$.
@@
$E[Y|D,X,Z]$ $=D\alpha_{1}+X^{\prime}\alpha_{2}+\left(1+\sum_{s=1}^{k_{g}}(\theta_{s}^{d}D^{s}+X^{\prime s}\theta_{s}^{x})\right)\left(\sum_{j=1}^{k_{v}}\mu_{j}\tilde{V}^{j}\right),$ (6)
and let $W$ denote all the regressors, but $D$. Under Assumptions 15, then $\alpha_{1}$ is identified as the coefficient of $D$ in the above regression and is given by
$\alpha_{1}$ $=\frac{E[(D-L_{D|W}(W))Y]}{E[(D-L_{D|W}(W))^{2}]},$ (7)
provided $E[(D-L_{D|W}(W))^{2}]\neq 0$.
4 Estimation
This section proposes a practical estimator for the parameter of interest $\alpha_{1}$, and establishes its statistical limiting properties.
4.1 An augmented control function estimator
We consider an estimator that consists of a two-step procedure. The overall procedure is as following. In the first-stage, one estimates the control function (CF), $\tilde{V}$. In the second-stage, one performs a simple OLS regression of the dependent variable $Y$ on the regressors $(D,X^{\prime})^{\prime}$ and the estimated CF, $\widehat{V}$, as well as on their polynomial interactions.
The goal is to estimate $\alpha_{1}$ given in equation (7). Recall that from the identification results discussed in the previous section, the population parameter is given by
$\alpha_{1}=\frac{E[(D-L_{D|W}(W))Y]}{E[(D-L_{D|W}(W))^{2}]},$
@@ 4.1 An augmented control function estimator
Denote the preliminary estimator of $\tilde{V}$ denoted by $\widehat{V}$. The estimator is given by
@@
where $\widehat{L}[D_{i}|\widehat{W}_{i}]=\widehat{W}_{i}^{\prime}(\widehat{W}^{\prime}\widehat{W})^{-1}\widehat{W}^{\prime}D.$ Here $\widehat{W}_{i}$ denotes the vector $W_{i}$ with $\tilde{V}_{i}$ replaced by $\widehat{V}_{i}$.
To implement this strategy in practice, we first estimate the errors in the first-stage regression assuming a linear heteroskedastic model.
@@
Assume $D=m(X,Z,V)=Z^{\prime}\pi_{1}+X^{\prime}\pi 2+h(X,Z)V$, with $E[V^{2}|X,Z]=1$ and $|h(X,Z)|>0$ almost surely.
The CF would be then the homoskedastic transformation of the first-stage residual, i.e. $\tilde{V}=V$ as in Assumption 6. To this end, we employ OLS in the first-stage and regress the endogenous variable $D$ on the instruments, $Z$, and exogenous variables $X$, to compute the residuals
Next, we have to compute the skedastic function $h(\cdot)$. Once we have an estimate of skedastic function, $\widehat{h}$, we obtain the normalized errors as
$\widehat{V}_{i}=\check{V}_{i}/\widehat{h}_{i},$
where $\widehat{h}_{i}=h(X_{i},Z_{i})$ is the estimated skedastic function.
The following assumption imposes a standard regularity condition on the skedastic function determining the smoothness and existence of moments of the conditional heteroskedasticity.
Assumption 7.
Assume that $h(X,Z)=h(X,Z;\gamma)$ a parametric function with parameter $\gamma$. (i) The map $\gamma\mapsto h(x,z;\gamma)$ is twice differentiable with gradient denoted by $\nabla_{\gamma}h(x,z;\gamma)$; (ii) $E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]$ has full rank; and (iii) the random vectors $Z_{i}h(X_{i},Z_{i})V_{i}$, $X_{i}h(X_{i},Z_{i})V_{i}$ and $\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}$ have finite second moments.
In practice, to estimate the skedastic function we employ the parametric models in Romano and Wolf (2017) or Wooldridge (2012, ch.8). For instance,
$h(X,Z;\gamma)^{2}=|Z^{\prime}|\gamma_{1}+|X^{\prime}|\gamma_{2}.$ (8)
or
$h(X,Z;\gamma)^{2}=\exp\left(\log|Z^{\prime}|\gamma_{1}+\log|X^{\prime}|\gamma_{2}\right).$ (9)
The coefficients of these models can be estimated using (possibly non-linear) OLS regressions with $\check{V}^{2}_{i}$ as the dependent variable. The resulting estimator is $\widehat{h}_{i}=h(X_{i},Z_{i};\widehat{\gamma})$.
Finally, in the second-step a standard OLS estimation is performed by regressing the dependent variable $Y$ on the regressors $(D,X^{\prime})^{\prime}$, the estimated CF, $\widehat{V}$, and their polynomial interactions, $\widehat{W}$, as in (6).
4.2 Asymptotic properties
From Assumptions 6 and 7, $V$ is given by
$V(\phi)$ $:=\frac{D-Z^{\prime}\pi_{1}-X^{\prime}\pi_{2}}{h(X,Z;\gamma)},$ (10)
where $\phi:=(\pi_{1}^{\prime},\pi_{2}^{\prime},\gamma^{\prime})^{\prime}$. For notational brevity, we write equation (6) as
$Y=D\alpha_{1}+W(\phi)^{\prime}\alpha_{w}+U,$ (11)
where $W(\phi)$ contains all the regressors except $D$, and by construction, $E[U|D,W]=0$.
The next result derive the influence function of $\widehat{\phi}$ and to derive the asymptotic properties of $\alpha_{1}$, the parameter of interest.
Proposition 1.
Under Assumptions 1, 2, 6, and 7, $\widehat{\phi}$ is asymptotically normal with influence function given by
$\sqrt{n}(\widehat{\phi}-\phi)$ $=\Sigma_{\phi}^{-1}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}M_{i}+o_{p}(1),$

### v4

@@ 3 A control function approach
Control Function Condition: $E(\varepsilon|D,X,Z)=E(\varepsilon|V)=r(V)$.
Assumption 3 is a standard control function condition, see, e.g., BlundellPowell08. This is the same condition as Assumption A-3 in FlorensHeckmanMeghirVytlacil08. Under this condition, the conditional expectation of the response variable in equation (1) can be written as
@@
$=D\alpha_{1}+X^{\prime}\alpha_{2}+g(D,X)r(V).$ (4)
It should be noted that, from the previous display, the skedastic term in the structural equation does not depend only on the control function, but also on $g$. Further assumptions are required to identify $\alpha_{1}$. To illustrate this, we use Example 2, where we assume that both $D$ and $Z$ are continuous.
@@
Consider the model in Example 1 without covariates, and assume Assumption 3 holds. Then (3) becomes $E[Y|D,Z]=D\alpha_{1}+g(D)r(\psi(D,Z))$, where $V:=\psi(D,Z)$ is identified from the first stage. Define $f(D,Z):=E[Y|D,Z]$. If $D$ and $Z$ are continuous, consider the following derivatives
$\frac{\partial f}{\partial D}$ $=\alpha_{1}+g_{D}r+gr_{V}\psi_{D}$
$\frac{\partial f}{\partial Z}$ $=gr_{V}\psi_{Z},$
@@
Note that, even though $\psi_{Z}$ and $\psi_{D}$ are identified, one cannot identify $\alpha_{1}$ because of the additional term $g_{D}r$, that is,
@@
The previous example shows that further assumptions are required for identification of the parameter $\alpha_{1}$. Assumption 4 below imposes a polynomial function on the skedastic function.
@@
This condition is the same as in FlorensHeckmanMeghirVytlacil08, except that here it includes additional exogenous covariates. [^4] FlorensHeckmanMeghirVytlacil08 write Assumption 4 directly in their model in equations (1) and (2). In Assumption 4, $X^{\prime k_{g}}$ refers to the element-by-element $k_{g}$ power of $X^{\prime}$. This could be replaced by any other polynomial that may include interactions.
Finally, we impose the following polynomial structure on the CF.
@@
Polynomial Control Function: $r(V)$ is a polynomial of degree $k_{v}$ in $V$ that does not contain a constant. That is $r(V)=\sum_{j=1}^{k_{v}}\mu_{j}{V}^{j}$.
This assumption is a trade-off between flexible parameterization and the nonparametric model used in (FlorensHeckmanMeghirVytlacil08, p2202). Finally, the identification results below also require a full rank condition that implies that the functions in Assumptions 4 and 5 are not linearly dependent with the full set of covariates $(D,X)$. Assumption 5 achieves this by not including a constant term in the function $r(V)$, but different conditions can be adapted as long as the full rank is satisfied, i.e. we may allow for $r(V)$ to include a constant, but restrict $g(D,X)$ to not contain a linear term with $D$ and $X$. The intuition for these restrictions is simple: we cannot have the same covariate to appear both in the linear function and in the implied control function expanded with the skedastic function. Suppose that there is only one endogenous regressor $D$ and that both the skedastic function and the CF contain a constant as: $g(D)=1+\delta D$ and $r(V)=1+V$. Then, $g(D)r(V)=1+V+\delta D+\delta DV$, and in this case, one is not able to identify $\alpha_{1}$ because of the term $\delta D$.
Identification of the parameter of interest requires identification of $V$, which in turn requires identification of $h$, which we can achieved up-to-scale. See for instance, RomanoWolf17 and Wooldridge12 for parametric specifications, and JiSuXiao15 and LintonXiao19 for nonparametric specifications. The next assumption imposes a normalization on the second moment of the control function, and requires the skedastic function to be positive.
Assumption 6.
Either $h\equiv 1$ or $E[V^{2}|X,Z]=1$ and $h(X,Z)>0$ almost surely.
The next lemma considers the identification of $\alpha_{1}$. It shows that the parameter of interest is identified by a regression of $Y$ on transformations of $(D,X,V)$. Define the linear projection of $D$ onto $W$ at $w$ as $L_{[D|W]}(w):=w^{\prime}E[WW^{\prime}]^{-1}E[WD]$, where $W$ denotes all the regressors in equation (2) except $D$, i.e. this contains $X^{\prime}$ and the interactions of $\{1,D^{s},X^{{}^{\prime}s}\}_{s=1,...,k_{g}}$ with $\{{V}^{j}\}_{j=1,...,k_{v}}$.
@@
$E[Y|D,X,Z]$ $=D\alpha_{1}+X^{\prime}\alpha_{2}+\left(1+\sum_{s=1}^{k_{g}}(\theta_{s}^{d}D^{s}+X^{\prime s}\theta_{s}^{x})\right)\left(\sum_{j=1}^{k_{v}}\mu_{j}{V}^{j}\right)$
$=D\alpha_{1}+W^{\prime}\alpha_{w}.$ (6)
Under Assumptions 16, $\alpha_{1}$ is the coefficient of $D$ in (2) and is given by
$\alpha_{1}$ $=\frac{E[(D-L_{[D|W]}(W))Y]}{E[(D-L_{[D|W]}(W))^{2}]},$ (7)
provided $E[(D-L_{[D|W]}(W))^{2}]\neq 0$.
Note that if $D$ is a dummy variable, as in the empirical application, then there is no need to consider a polynomial model of it and the regression is simplified. The identification result in Lemma 2 above shows that the parameter of interest $\alpha_{1}$ in equation (7) can be written as a function of observable data. In the next section, we propose a simple two-step estimator for it.
4 Estimation and inference
This section proposes a practical estimator for the parameter of interest $\alpha_{1}$, and establishes its statistical limiting properties. We consider an estimator that consists of a two-step procedure. The overall procedure is as following. In the first-stage, one estimates the control function (CF), $V$. In the second-stage, one performs a simple OLS regression of the dependent variable $Y$ on the regressors $(D,X^{\prime})^{\prime}$ and the estimated CF, $\widehat{V}$, as well as on their polynomial interactions. In addition, we provide practical inference procedures.
4.1 First stage estimator: obtaining $\hat{V}$
We employ OLS in the first-stage and regress the endogenous variable $D$ on the instruments, $Z$, and exogenous variables $X$, to compute the residuals
Next, we compute the skedastic function $h(\cdot)$. In practice, to estimate the skedastic function we employ the parametric models in RomanoWolf17 or Wooldridge12. For instance,
$h(X,Z;\gamma)=\left(|Z^{\prime}|\gamma_{1}+|X^{\prime}|\gamma_{2}\right)^{1/2}.$ (8)
or
$h(X,Z;\gamma)=\left(\exp\left(\log|Z^{\prime}|\gamma_{1}+\log|X^{\prime}|\gamma_{2}\right)\right)^{1/2}.$ (9)
[…]
$\widehat{V}_{i}=\check{V}_{i}/\widehat{h}_{i}.$
The following assumption imposes a standard regularity condition on the skedastic function determining its smoothness and existence of moments of the conditional heteroskedasticity.
Assumption 7.
Assume that $h(X,Z)=h(X,Z;\gamma)$, a parametric function with parameter $\gamma$. (i) The map $\gamma\mapsto h(x,z;\gamma)$ is twice differentiable with gradient denoted by $\nabla_{\gamma}h(x,z;\gamma)$; (ii) $E[\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}]$ has full rank; (iii) the vectors $Z_{i}$, $X_{i}$ and the components of $\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\nabla_{\gamma}h(X_{i},Z_{i};\gamma)^{\prime}h(X_{i},Z_{i};\gamma)^{2}$ have finite second moments; and (iv) the vectors $Z_{i}h(X_{i},Z_{i})V_{i}$, $X_{i}h(X_{i},Z_{i})V_{i}$ and $\nabla_{\gamma}h(X_{i},Z_{i};\gamma)\left(V_{i}^{2}-1\right)h(X_{i},Z_{i};\gamma)^{3}$ have finite second moments.
Denote $\phi:=(\pi_{1}^{\prime},\pi_{2}^{\prime},\gamma^{\prime})^{\prime}$. The next result is the asymptotic normality of $\widehat{\phi}=(\hat{\pi}_{1}^{\prime},\hat{\pi}_{2}^{\prime},\hat{\gamma}^{\prime})^{\prime}$. The precise form of the influence function $\psi_{\phi,i}$ is detailed in the appendix.
Proposition 1.
Under Assumptions 1, 2, 6, and 7,
$\sqrt{n}(\widehat{\phi}-\phi)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\phi,i}+o_{p}(1)\overset{d}{\to}\mathcal{N}(0,\Omega_{\phi}),$
where $\Omega_{\phi}=E[\psi_{\phi}\psi_{\phi}^{\prime}]$.
4.2 Second stage estimator: augmented control function estimator
In the second step, the goal is to estimate $\alpha_{1}$ given in equation (7). Recall that from the identification results discussed in the previous section, the population parameter is given by
$\alpha_{1}=\frac{E[(D-L_{[D|W]}(W))Y]}{E[(D-L_{[D|W]}(W))^{2}]},$
@@ 4.2 Second stage estimator: augmented control function estimator
Denote the preliminary estimator of $V$ denoted by $\widehat{V}$. The estimator is given by
@@
where $\widehat{L}[D|\widehat{W}_{i}]=\widehat{W}_{i}^{\prime}(\widehat{W}^{\prime}\widehat{W})^{-1}\widehat{W}^{\prime}D.$ Here $\widehat{W}_{i}$ denotes the vector $W_{i}$ with $V_{i}$ replaced by $\widehat{V}_{i}$.
Given that the standard CF model (i.e. adding $\check{V}$, the first-stage residuals, to the regression equation) is in fact the 2SLS estimator, this determines that the latter is a special case of our augmented CF procedure. That is, 2SLS-IV estimator is nested in our model. The presence of heteroskedasticity in the structural model and its subsequent effect can be evaluated by adding the interaction term $\check{V}D$ and checking its statistical significance. If the interaction is not statistically significant, this provides some evidence that the 2SLS might not be biased. This has a similar intuition to a test for endogeneity in linear regression models when checking the statistical significance of $\check{V}$ in the CF-2SLS model (see the examples in Wooldridge10, ch.6). As such, our proposed method is robust to the presence of homoskedasticity and heteroskedasticity in the structural equation. However, the inclusion of additional terms and the previous estimation of the CF in the reduced form OLS estimation might impact the efficiency of the augmented CF procedure.
For notational brevity, we write equation (2) as
$Y=D\alpha_{1}+W(\phi)^{\prime}\alpha_{w}+U,$ (10)
where $W(\phi)$ contains all the regressors except $D$ as given in Lemma 2, and by construction, $E[U|D,W]=0$. Furthermore, $\alpha=(\alpha_{1},\alpha_{w}^{\prime})^{\prime}$. The next result is the asymptotic normality of $\widehat{\phi}=(\hat{\pi}_{1}^{\prime},\hat{\pi}_{2}^{\prime},\hat{\gamma}^{\prime})^{\prime}$.
The next theorem contains the limiting distribution of the proposed two-step CF estimator of $\alpha=(\alpha_{1},\alpha_{w}^{\prime})^{\prime}$, denoted $\widehat{\alpha}=(\widehat{\alpha}_{1},\widehat{\alpha}_{w}^{\prime})^{\prime}$. This provides details for the asymptotics for the estimator of interest, i.e., $\alpha_{1}$, which is an application of OLS with generated regressors. We consider the following assumption:
@@
Let $W(\phi)$ be all the regressors in (10) (except $D$) as given in Lemma 2, and define $R_{i}(\phi):=(D_{i},W_{i}(\phi)^{\prime})^{\prime}$. Assume that uniformly in $\phi$ (i) $E[R_{i}(\phi)R_{i}^{\prime}(\phi)]$ is non-singular; (ii) a (uniform) law of large numbers holds for $\frac{1}{n}\sum_{i=1}^{n}R_{i}(\phi)R_{i}^{\prime}(\phi)$, $\frac{1}{n}\sum_{i=1}^{n}(U_{i}J_{\phi,i}-R_{i}(\phi)\alpha^{\prime}J_{\phi,i})$, where $J_{\phi}$ is the Jacobian of $\phi\mapsto R(\phi)$, $J_{\phi}=\nabla_{\phi}R(\phi)$; and (iii) $E[\|RR^{\prime}U^{2}\|]<\infty$.
Assumption 8 $(i)$ ensures that the CF additional terms to the structural equation are not perfectly collinear to the main regressors, and Assumptions 8 $(ii)$ and $(iii)$ are required for the asymptotic representations using the influence functions. [^5] Sufficient conditions for a uniform law of large numbers can be found in Lemma 2.4 in NeweyMcFadden1994. Essentially, it depends on a dominating function for $R(\phi)$ or, alternatively, $h(x,z;\gamma)$.
Theorem 1.
Under Assumptions 1-8,
$\sqrt{n}(\widehat{\alpha}-\alpha)$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{\alpha,i}+o_{p}(1)\overset{d}{\to}\mathcal{N}(0,\Omega_{\alpha})$
where $\psi_{\alpha,i}:=E[RR^{\prime}]^{-1}R_{i}U_{i}+E[RR^{\prime}]^{-1}E[UJ_{\phi}-R(\phi)\alpha^{\prime}J_{\phi}]\psi_{\phi,i}$, and $\Omega_{\alpha}=E[\psi_{\alpha}\psi_{\alpha}^{\prime}].$
Note that $\psi_{\phi,i}$ is contributing to $\psi_{\alpha,i}$, the influence function of $\hat{\alpha}$. From the above result, we obtain $\sqrt{n}(\widehat{\alpha}_{1}-\alpha_{1})\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\Omega_{\alpha,(1,1)}),$ where $\Omega_{\alpha,(1,1)}$ is the $(1,1)$ element of $\Omega_{\alpha}$.
4.3 Inference
Now we turn our attention to inference in the model under endogenous heteroskedasticity, and suggest a practical estimator for the asymptotic variance-covariance matrix. We also formally establish its consistency.
Estimation of $\Omega_{\alpha,(1,1)}$ can be carried out by resorting to the sample counterpart of $\Omega_{\alpha}$ as following:
$\widehat{\Omega}_{\alpha}$ $=\frac{1}{n}\sum_{i=1}^{n}\hat{\psi}_{\alpha,i}\hat{\psi}_{\alpha,i}^{\prime}.$
with
$\hat{\psi}_{\alpha,i}$ $=\left(\frac{1}{n}\sum_{j=1}^{n}\widehat{R}_{j}\widehat{R}_{j}^{\prime}\right)^{-1}\widehat{R}_{i}\widehat{U}_{i}+\left(\frac{1}{n}\sum_{j=1}^{n}\widehat{R}_{j}\widehat{R}_{j}^{\prime}\right)^{-1}\left(\frac{1}{n}\sum_{j=1}^{n}\left(\widehat{U}_{j}J_{\hat{\phi},j}-\widehat{R}_{j}\widehat{\alpha}^{\prime}J_{\hat{\phi},j}\right)\right)\hat{\psi}_{\phi,i},$